A disk between vertebrae in the spine is subjected to a shearing force of 690 N. Find its shear deformation taking it to have the shear modulus of 1.00 109 N/m2. The disk is equivalent to a solid cylinder 0.700 cm high and 4.00 cm in diameter.

_______m?
Answer #1

A shearing force of Fs=690 N is acting on the disk between vertebrae in the spine.

The disk is equivalent to a solid cylinder L=0.700 cm=0.7×10−2 m high and d = 4.00 cm = 4.00×10−2 m in diameter.

The shear modulus of the disk is G=1.00×109 N/m2.

Shear area is given by

Shear strain for a lateral deformation of Δs is given by

From the expression of Shear modulus we have